A rectangular swimming pool measures 20 feet by 40 feet. A walk
of uniform width surrounds it. The total area of the pool and the
walk is 1500 ft². Find the width of the walk.

Answer by mananth(16946)

width = x
length of pool + walk = 40+2x
with of pool + walk = 20+2x
Total area = (40+2x)(20+2x)= 1500
800+80x+40x+4x^2=1500
4x^2+120x+800-1500=0
4x^2+120x-700=0
/4
x^2+30x-175=0
x^2+35x-5x-175=0
x(x+35)-5(x+35)=0
(x+35)(x-5)=0
x= 5 the width. x cannot be -35
m.ananth@hotmail.ca